OpenStudy (anonymous):

At Water World the admission fee is $17.00, and you can rent an inner tube for $0.75 per hour. You can use the formula P = 17.00 + 0.75h to find the total cost for admission plus renting an inner tube. a. Solve the formula P = 17.00 + 0.75h for h. Explain your steps for solving the formula for h. b. For how many hours can you rent an inner tube if you only have $20.75 to spend? Explain your method for finding the answer.

OpenStudy (bahrom7893):

P = 17.00+0.75h Subtract 17.00 from both sides: P - 17.00 = 0.75h Divide both sides by 0.75: (P-17.00)/0.75 = h

OpenStudy (bahrom7893):

b) If you have $20.75, then let's see how many hours we can get: h = (P-17.00)/0.75 = (20.75-17.00)/0.75 = 3.75/0.75

OpenStudy (bahrom7893):

h = 5 hrs

OpenStudy (anonymous):

thankkk youuuuu(:

OpenStudy (anonymous):

bahroom is good at this stuff

OpenStudy (bahrom7893):

no Ts or double Os